# A Novel Knapsack-based Financial Portfolio Optimization using Quantum Approximate Optimization Algorithm

### v2

@@ 4.3.3 Number of Qubit Requirements
The presented illustration implies that using a feasibility oracle with nf ancilla qubits requires $N+n_{f}+3$ qubits. In contrast, there is an exigency for an additional three qubits, specifically for generating the unitary $U_{B}$ (Equation 33).
@@ 5.1 Experimental Setup
Initially, we implemented the proposed method using the Qiskit library [37], a widely used open-source quantum computing framework. We run the algorithms on the QASM simulator, a tool provided by Qiskit for simulating quantum circuits, which helped us gain insights into the behavior and performance of the algorithms before moving on to the real quantum device. While using the real device, we used IBM Cairo, a 27-qubit quantum device, for quantum computations, which offers the opportunity to test the algorithms in a real-world quantum computing environment. The experimental phase involves conducting several tests in which we select varying stocks from well-known companies, such as Apple Inc. (AAPL), Amazon.com Inc. (AMZN), Alphabet Inc. (GOOGL), Microsoft Corporation (MSFT), NVIDIA Corporation (NVDA), and Tesla, Inc. (TSLA). Notably, the selection of these stocks was based on different scenarios, including cases with two, three, four, and five stocks. Postselection, we proceeded with the experimental setup of the proposed method. During the optimization process, we focused on optimizing QAOA, which was achieved by optimizing the $2p$ angle parameters, i.e., $\beta$ and $\gamma$, using the classical optimizer SHGO [38]. Additionally, we integrated quantum walk to boost optimization by leveraging its benefits for refining the process and achieving superior results. To evaluate the performance of the proposed method, we defined the problem of our algorithm based on the number of assets to be optimized (Table 1). First, we selected 2–5 subsets of stocks using the capabilities of PyPortfoliopt [39]. The values in Table 1 represent the expected return of stocks counting from January 1, 2018, to January 1, 2023. This step clarifies the problem and enables us to tailor our algorithm accordingly. We aimed to evaluate the effectiveness of the QAOA optimization technique concerning portfolio optimization. To achieve this, we specifically evaluated the performance of the algorithm using the best-known solution (BKS), a result of using a classical algorithm [40], where $1$ signifies selection and 00 denotes exclusion. This technique compared the results obtained from our algorithm with those of the classical solution, providing a benchmark for assessing its performance and effectiveness. By undertaking this evaluation, we can gain valuable insights into the capabilities and limitations of the QAOA optimization technique in portfolio optimization.
# Stocks Type Values Capacity BKS Qubit
2 MSFT 0.26930 1 (1,0) 7
AAPL 0.2488
3 MSFT 0.2315 1 (1,0,0) 8
AAPL 0.2208
NVDA 0.3638
4 MSFT 0.2089 2 (1,1,0,0) 10
AAPL 0.1984
GOOGL 0.2037
NVDA 0.3220
5 MSFT 0.1858 2 (0,1,0,0,1) 11
AAPL 0.1941
GOOGL 0.1777
AMZN 0.1826
NVDA 0.2834
Table 1: Preparation for the portfolio optimization model uses expected returns based on the Markowitz model, as values in a knapsack problem, along with a key constraint. Capacity, set at $1$ or $2$, limits the stocks to $1$ or $2$, respectively. The model includes a classical solution best-known solution for comparing performance and specifies the number of qubits used in each stock selection scenario.
@@ 5.2 Performance Evaluation
By meticulously exploring the variable settings within the QWM-QAOA, we identified a circuit layer of $p=5$ and a trotter step value of $m=5$ as the optimal configuration for achieving superior results. Contextually, we explore the probability distributions arising for each stock in the portfolio. Figure 5 shows the experimental results, which are the optimal results of each scenario of the number of stocks. The obtained results reveal the probability distribution results for all possible cases of stock selection. Additionally, we employed the BKS method to evaluate the obtained results and ensured its expected behavior (Table 1).
Within the circuit layer of $5$, the approximation ratio varied based on the number of stocks in the portfolio. The obtained results ranged from 100% to 98% in a noiseless environment, from 98% to 80% when the fake backend was used, and from 49% to 33% when the real device was employed. When we extended our study to real-world conditions by implementing the experiment on the IBM Cairo device, the approximation ratio decreased to approximately 50% because of gate errors and decoherence affecting the fidelity of quantum states [41]. Addressing these challenges will require additional efforts, such as error mitigation, which is beyond the scope of this study [42, 43].
Figure 5: Probabilities for different scenarios involving various numbers of stocks under three distinct environmental conditions: noiseless with the simulator, noisy with FakeCairoV2, and noisy with the ibm_cairo device.
@@ 5.2.1 Sensitivity Analysis
We conducted a comprehensive sensitivity analysis of key parameters using the QWM–QAOA algorithm. Our empirical exploration encompassed crucial elements, such as the circuit layer $(p)$ and the number of trotter steps $(m)$. These parameters are crucial for shaping the effectiveness of the proposed algorithm, and variation comprehension of their sensitivity is crucial for optimizing the performance of the algorithm. Through a methodical examination and evaluation, we present details of the relationship between these parameters and how much they influence the overall efficiency and resilience of QAOA.
  • Circuit layer: Figure 6 shows that as the circuit layer increases beyond the initial value of $p\geqslant 3$, the approximation ratio of the diverse portfolio optimization result consistently reaches an impressive result ranging from 100% to 98%, categorically based on the setup cases of portfolio optimization. This empirical evidence emphatically establishes that our performance consistently produces an efficient result even at some circuit layers.
[…]
  • Trotter step count in QWM-QAOA: We aimed to find the parameter $m$, representing the trotter step counts in the QWM–QAOA algorithm while maintaining a constant circuit layer of $p=3$ and optimize $\beta$, $\gamma$ for the resulting circuit of 11115555. Table 2 shows that the approximation ratio for various stock selections increasingly exhibits a better result starting from $m=3$, ranging from 99% to 97%. Finally, when $m=5$, we obtained efficient optimization results from our proposed method ranging from 100% to 98%.
@@
    2 1 1 0.99 1 1
    3 0.99 0.61 0.99 0.99 1
    4 0.99 0.99 0.99 0.99 0.99
    5 0.80 0.97 0.97 0.98 0.98
    Table 2: Number of trotters step $(m)$ of QWM–QAOA from $1$ to $5$, followed by the corresponding approximation ratio; results emphasize the importance of careful consideration when selecting the trotter step count, as it can significantly impact the approximation ratio.
6 Discussion and Conclusion
We empirically investigate the effectiveness of quantum computing techniques in portfolio optimization by transforming the problem into a knapsack-based format and using QWM–QAOA as the optimization method. Table 1 shows the relationship between the number of stocks and required qubits. We compare our results with conventional optimization (BKS) to ensure accuracy. Optimization was performed under three different scenarios (Figure 5), achieving efficient results in noiseless and noisy device settings (100% to 98% and 98% to 80%, respectively). However, in noisy real devices fidelity drops to around 50%, highlighting the need for further research in error mitigation. Circuit layer analysis (Figure  6) shows our method yields 100% to 98% accuracy with $p\geqslant 3$ for various scenarios of 2–5 stocks. Table 2 emphasizes the importance of Trotter steps $(m)$ selection in QWM–QAOA, with $m=5$ reaching 100% to 98% approximation. By extending experiments to larger stock configurations, we contribute insights into the proposed approach’s applicability and efficiency across a broader spectrum of portfolio sizes. This research enhances our understanding of quantum computing’s potential in portfolio optimization, particularly for the knapsack-based problem in the financial domain. Future work will focus on designing error mitigation techniques for larger datasets to further improve the method’s potential.
Acknowledgments
This work was supported by a Research Grant of Pukyong National University(2023).
Data availability statement
@@ Acknowledgments
The data that support the findings of this study are openly available at the following https://github.com/QCL-PKNU/QAOA-Knapsack-Portfolio-Optimization.git

### v3

@@ 4.3.3 Number of Qubit Requirements
The presented illustration implies that using a feasibility oracle with $n_{f}$ ancilla qubits requires $N+n_{f}+3$ qubits. In contrast, there is an exigency for an additional three qubits, specifically for generating the unitary $U_{B}$ (Equation 32).
@@ 5.1 Experimental Setup
We present the experimental setup for evaluation, specifically the datasets, hyperparameter configuration, and model setup used to evaluate our proposed approach.
5.1.1 Hyperparameter Configuration
@@
Initially, we implemented the proposed method using the Qiskit library [42], a widely used open-source quantum computing framework. We run the algorithms on the QASM simulator, a tool provided by Qiskit for simulating quantum circuits, which helped us gain insights into the behavior and performance of the algorithms before moving on to the real quantum device. Moreover, we used a fake backend noise from 127 qubits of IBM FakeWashington. While using the real device, we used IBM Nazca, a 127-qubit quantum device, for quantum computations, which offers the opportunity to test the algorithms in a real-world quantum computing environment. The experimental phase involves conducting several tests in which we select varying stocks from $yahoofinance$, such as Apple Inc. (AAPL), Amazon.com Inc. (AMZN), Alphabet Inc. (GOOGL), Microsoft Corporation (MSFT), NVIDIA Corporation (NVDA), and Tesla, Inc. (TSLA), Netflix, Inc. (NFLX), and Visa Inc. (N). Notably, the selection of these stocks was based on different scenarios, including all cases from two to eight stocks. Postselection, we proceeded with the experimental setup of the proposed method. During the optimization process, we focused on optimizing QAOA, which was achieved by optimizing the $2p$ angle parameters, i.e., $\beta$ and $\gamma$, using the classical optimizer SHGO [43]. Additionally, we integrated quantum walk to boost optimization by leveraging its benefits for refining the process and achieving superior results.
Table 1: Hyperparameter Configuration
Hyperparameter Value
Number of circuit layers 3
Number of trotter steps 3
Financial data provider yahoo finance
Portfolio valuation module PyPortfolioOpt
Classical optimizer SHGO
Framework for architecture implementation Qiskit
5.1.2 Problem Setup
To evaluate our proposed method’s performance, we defined our algorithm’s problem based on the number of assets to be optimized (Table 2). First, we selected 2–8 subsets of stocks using the capabilities of PyPortfoliopt [44]. The values represent the expected return of stocks counting from January 1, 2018, to January 1, 2023. The knapsack capacity $Cap$ is the maximum weight, where in our case we set it to be half of the total number of assets. This step clarifies the problem and enables us to tailor our algorithm accordingly. We aimed to evaluate the effectiveness of the QAOA optimization technique in portfolio optimization. To achieve this, we specifically evaluated the performance of the algorithm using the best-known solution (BKS), a result of using a classical algorithm [45], where $1$ signifies selection and 00 denotes exclusion. This technique compared the results obtained from our algorithm with those of the classical solution, providing a benchmark for assessing its performance and effectiveness. By undertaking this evaluation, we can gain valuable insights into the capabilities and limitations of the QAOA optimization technique in portfolio optimization.
Table 2: The optimal stock selection and values for various portfolio sizes under the binary knapsack constraint. It includes the number of selected stocks (#Ss), stock types (Type), their values, knapsack capacity (Cap), Best Known Solution (BKS), and qubits used (#Qs). The BKS column shows the binary string representation of the optimal solution.
#Ss Type Values Cap BKS #Qs
2 MSFT 0.2430 1 (0,1) 7
AAPL 0.2602
3 MSFT 0.2430 1 (0,1,0) 8
AAPL 0.2602
NVDA 0.2430
4 MSFT 0.2430 2 (1,1,0,0) 10
AAPL 0.2602
GOOGL 0.1047
NVDA 0.2430
5 MSFT 0.2430 2 (0,1,0,0,1) 11
AAPL 0.2602
GOOGL 0.1047
AMZN 0.0716
NVDA 0.2430
6 MSFT 0.2430 3 (0,1,0,0,1,1) 12
AAPL 0.2602
GOOGL 0.1047
AMZN 0.0716
NVDA 0.2430
TSLA 0.4203
7 MSFT 0.2430 3 (0,1,0,0,1,1,0) 13
AAPL 0.1870
GOOGL 0.1749
AMZN 0.1898
NVDA 0.2856
TSLA 0.4203
NFLX 0.0797
8 MSFT 0.2430 4 (1,1,0,0,1,1,0,0) 15
AAPL 0.1899
GOOGL 0.1780
AMZN 0.1903
NVDA 0.2874
TSLA 0.4203
NFLX 0.0797
V 0.1341
@@ 5.2 Performance Evaluation
In this section, we present our model performance and its sensitivity analysis.
5.2.1 Model Performance
By meticulously exploring the variable settings within our QWM-QAOA, we identified that a circuit layer depth of $p=3$ and a Trotter step value of $m=3$ yielded the best results. Figure 5 illustrates the approximation ratios for knapsack-based portfolio optimization with varying numbers of stocks (ranging from 2 to 8 stocks) under three different environmental conditions: noise-free (simulator), noisy (using FakeWashington as a simulated backend), and noisy (real device: IBM Nazca with 127 qubits).
  • Noise-free (Simulator): The approximation ratios remain very high across all stock configurations, ranging from 100% (for 2 and 3 stocks) to 96% (8 stocks), indicating that the algorithm performs near-perfectly in an ideal, noise-free environment.
  • Noisy (Fake Backend): The introduction of noise through the simulated FakeWashington backend results in a decline in performance. The approximation ratios range from 98% to 70%, with lower values observed for larger portfolios. For example, 2 stocks achieve an approximation ratio of 98%, while 6 and 7 stocks drop to 70%. Interestingly, for 8 stocks, the approximation ratio slightly improves to 78%, indicating a non-linear impact of noise for larger portfolios.
  • Noisy (Real Device): The experiment is executed on the real IBM Nazca device with the approximation ratios dropping further, ranging from 96% (for 2 stocks) to 51% (for 6 stocks). Surprisingly, for 7 and 8 stocks, the approximation ratios increased to 52% and 55%, respectively.
This performance reduction is due to real-world factors, such as gate errors and decoherence, which exert a more significant impact as the number of stocks increases. This unexpected result hints at the potential for further improvements, even as the number of stocks increases in a noisy environment [46]. These issues highlight the importance of advanced error mitigation techniques, though such methods are beyond the scope of this study [47, 48].
Figure 5: Approximation ratio for different scenarios involving various numbers of stocks under three distinct environmental conditions: noise-free with the simulator, noisy with Fake Washington, and noisy with the IBM Nazca device.
@@ 5.2.2 Sensitivity Analysis
We conducted a sensitivity analysis of key parameters using our QWM–QAOA algorithms, such as the circuit layer $(p)$ and the number of trotter steps $(m)$. These parameters are crucial for shaping the effectiveness of our proposed algorithm. Through a methodical examination and evaluation, we present details of the relationship between these parameters and how much they influence the overall efficiency and resilience of our QWM–QAOA.
  • Circuit Layer: Figure 6 shows that as the circuit layer increases beyond the initial value of $p\geqslant 3$, the approximation ratio of the diverse portfolio optimization result consistently reaches an impressive result ranging from 100% to 95%, categorically based on the setup cases of portfolio optimization. This empirical evidence emphatically establishes that our performance consistently produces an efficient result even at some circuit layers.
    Figure 6: Approximation ratio of the circuit layer $(p)$ is 1–5 across all problems. Remarkably, when $p\geqslant 3$, the approximation ratio reaches 100%, indicating that the proposed method is efficient even at the lower circuit layer.
  • Trotter Step in QWM-QAOA: We aimed to find the parameter $m$, representing the trotter step counts in the QWM–QAOA algorithm while maintaining a constant circuit layer of $p=3$ and optimize $\beta$, $\gamma$ for the resulting circuit of 1–5. Table 3 shows that the approximation ratio for various stock selections increasingly exhibits a better result starting from $m=3$, ranging from 100% to 95%. When $m=5$, we obtained efficient optimization results from our proposed method ranging from 100% to 96%. The approximation ratios demonstrate how increasing the trotter steps improves the portfolio optimization results, particularly when $m\geqslant 3$, yielding consistently higher performance across different portfolio problem sizes.
    Table 3: Performance of QWM-QAOA with varying Trotter steps across different stock selection scenarios. The table presents the approximation ratios for various numbers of stocks, ranging from 2 to 8, using the QWM-QAOA algorithm with different Trotter step values (1 to 5).
@@
    2 1.00 0.96 1.00 1.00 1.00
    3 1.00 0.60 1.00 1.00 1.00
    4 0.77 0.99 0.99 0.99 0.99
    5 0.96 0.96 0.98 0.98 0.98
    6 0.96 0.97 0.98 0.98 0.98
    7 0.96 0.96 0.95 0.96 0.96
    8 0.96 0.96 0.96 0.96 0.96
6 Discussion and Limitation
our research provides a novel application of quantum computing, specifically through the Quantum Walk Mixer combined with a Quantum Approximate Optimization Algorithm (QWM-QAOA) integrating with Quantum Fourier Transform (QFT), to address the knapsack-based financial portfolio optimization problem. By transforming the traditional portfolio optimization problem into a binary knapsack problem formulation, we were able to leverage quantum computing to provide efficient and effective optimization strategies, particularly for NP-hard problems. The results from our experiment on noise-free simulation indicate that the QWM-QAOA approach can yield high approximation ratios in both fake backend and real quantum devices, demonstrating the potential of quantum computing in financial optimization tasks. Our study also highlights the importance of various QAOA parameters, such as circuit layer $p$ and trotter step $m$, in optimizing the performance of our proposed algorithm. Through sensitivity analysis, we found that a circuit layer of $p\geqslant 3$ and the number of trotter steps $m\geqslant 3$ provided the most converged results, with approximation ratios consistently reaching between 100% and 96% in the best cases. This empirical evidence supports the notion that QAOA with the proper configuration, can outperform classical optimization approaches in the context of financial portfolio optimization problems. However, our experiments on real quantum hardware revealed some limitations. The approximation ratio dropped significantly due to unavoidable hardware noise such as gate errors and decoherence errors, with real-device performance yielding ratios as low as 50%. Despite these current limitations of quantum hardware, this finding emphasizes the pressing need for continued research in quantum error mitigation (QEM) techniques and hardware advancements to unlock the full potential of our proposed QWM-QAOA method on portfolio optimization problems in real-world applications.
Noise and Hardware Constraints: The most significant limitation encountered in this study was the impact of overall gate errors when using real quantum devices [49]. While the QAOA algorithm performed well in simulators and fake backend environments, real hardware results showed a substantial drop in approximation ratios (below 50%). This limitation reflects the current state of NISQ devices, which are not yet fully capable of handling complex optimization problems without significant fidelity loss.
Scalability: While our approach worked well for smaller portfolios, scaling QWM-QAOA to larger datasets faces challenges due to the increasing number of required qubits and the limitations of current quantum hardware, such as limited qubit availability and short coherence times. These constraints hinder its applicability to larger portfolio optimization problems. Advancements in quantum hardware and quantum error mitigation (QEM) techniques along with hybrid approaches are needed to extend QWM-QAOA to more complex real-world scenarios.
Limited Dataset: The experimental evaluations were conducted on a small selection of stocks, ranging from two to eight, providing a controlled setting to demonstrate the algorithm’s effectiveness. However, to determine its practical value, future evaluations must incorporate historical data, transaction costs, and market dynamics. Expanding the tests to larger and more diverse datasets will be crucial for assessing the broader applicability of the QWM-QAOA approach in more complex, real-world financial scenarios.
7 Conclusion
In this research, we proposed a novel approach to financial portfolio optimization by leveraging the capabilities of quantum computing, specifically through the Quantum Walk Mixer integrated with a Quantum Approximate Optimization Algorithm (QWM-QAOA). We reformulated the traditional portfolio optimization problem into a knapsack problem, enabling us to address the NP-hard nature of this financial decision-making challenge. Our experimental evaluations, conducted across simulated, fake, and real quantum devices, demonstrated the effectiveness of our approach in yielding high approximation ratios, particularly in noise-free environments. The results of our study reveal that the QWM-QAOA approach is capable of producing efficient portfolio selections, with approximation ratios ranging from 100% to 95% in noise-free conditions and from 98% to 70% on noisy fake devices. However, when tested on real quantum hardware, the results showed a decline in performance due to noise and gate errors, reducing the approximation ratios to approximately 50%. These findings highlight the current limitations of quantum hardware, particularly in dealing with noisy intermediate-scale quantum (NISQ) devices. While our proposed method showed promise, especially in noise-free environments, future work is needed to address the noise in real quantum devices. Error mitigation techniques with further exploration of quantum algorithms for optimization will be crucial in enhancing the robustness and scalability of this approach in real-world applications. Our study contributes to the growing field of quantum finance, offering insights into how quantum algorithms can be applied to complex financial problems such as portfolio optimization.
Code Availability
@@ 7 Conclusion
The code that supports the findings of this study is openly available in the Github repository, QWM-QAOA.
Acknowledgment
This research was supported by the "Regional Innovation Strategy (RIS)" through the National Research Foundation of Korea (NRF), funded by the Ministry of Education (MOE) (2023RIS-007).